A 1-in, constant diameter shaft is loaded with forces at A and B as shown, with ground reaction forces at O and C. The shaftDetermine the minimum static factor of safety using the distortion-energy failure theory. The minimum factor of safety is
Answer #1

46016 5751bf Izin isim lom AC Rc Ro To fond sealtics support 115 -O Rotre 575-460 = Efeco Emoco £60x12 +RC X 40 =575 X 30 575

SFD and Bm D SED 281075lb 0 B C A -178.25lb. 293.25 lb BMD 2139-1b-in bf-in 29 32.5 (Ь (87) max ax =281-75 Ibris com) 29 32.5

# Ibris T= 1500 6 167 Te + shear stress due to torsing! load Tid 3 Tau IT = 16x1500 7x13 = 7639044 1bf ho2 a direct Cdirect s

(i) 92 = obe - 1 (0)24 772 29875.8 (8012. 88) 士 2 31889.22 psi 11 얼 -2013.42 psil According to shear store so therry maximum